Each year, my student’s Mike’s kazoo collection appreciates (goes up in value) 10%. Mike says that after 10 years, the value will double. How do you respond?
Answer by stanbon(75887)

Each year, my student’s Mike’s kazoo collection appreciates (goes up in value) 10%. Mike says that after 10 years, the value will double. How do you respond?
----------
Assume the value now is "x".
After 1 year year the value is x + 0.1x = 1.1x
After the 2nd year the value is 1.1x + 0.1(1.1x) = 1.1^2x
After 10 years the value will be (1.1^10) x
=============
Now 1.1^10= 2.59...
So the value in 10 years will be (2.59..)x which is more that doubling the x.
=======================
Cheers,
Stan H.